OpenStudy (anonymous):

I need to look for the horizontal asymptote of (sin x)/(1 + cos x)...How I do this? :/

OpenStudy (anonymous):

cosx=-1 x=pie

OpenStudy (anonymous):

there are undefined when 1+cosx=0 this is horyzontal asymptote

OpenStudy (anonymous):

Do you know how can I resolve this using the limit x->infinity equation?

OpenStudy (anonymous):

this is what I'm talking about http://www.wolframalpha.com/input/?i=lim+x-%3Einf+%28sen%28x%29%29%2F%281%2B+cos%28x%29%29

OpenStudy (anonymous):

I need to find the asymptote using definition of limit when x-> infinity

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{sinx}{1+cosx}*\frac{1-cosx}{1-cosx}=\lim_{x \rightarrow \infty}\frac{sinx*(1-cosx)}{1-\cos^2x}\] \[=\lim_{x \rightarrow \infty}\frac{sinx(1-cosx)}{\sin^2x}=\lim_{x \rightarrow \infty}\frac{1-cosx}{sinx}=\lim_{x \rightarrow \infty}(cscx-cotx)\] there is no horizontal asymptote

myininaya (myininaya):

you can find the vertical asymptotes by solving this for x: 1+cosx=0

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

do you know any other method of how I can find horizontal asymptotes of (x+1)^3 / (x-1)^2 ?? without using L'Hospital

OpenStudy (anonymous):

using lim x->infinity....but not using L'Hospital

myininaya (myininaya):

ok do you remember \[\lim_{x \rightarrow \infty}\frac{1}{x^r}=0, r \in \mathbb{Z}\]

OpenStudy (anonymous):

yes

myininaya (myininaya):

\[\lim_{x \rightarrow \infty }\frac{(x+1)^3}{(x-1)^2}=\lim_{x \rightarrow \infty}\frac{x^3+3x^2+3x+1}{x^2-2x+1}\] now divide top and bottom by x^2 \[\lim_{x \rightarrow \infty}\frac{x+3+\frac{3}{x}+\frac{1}{x^2}}{1-\frac{2}{x}+\frac{1}{x^2}}\]

myininaya (myininaya):

top goes to infinity bottom goes to 1 so this means there is no horizontal asymptote

myininaya (myininaya):

now we didnt have to multiply that junk out i just thought it would be easier to see

myininaya (myininaya):

if you have polynomial/polynomial degree of top is equal to degree of bottom the horizontal asymptote is y=coefficient of the leading term (highest exponent) on top over coefficient of the leading term (highest exponent) on bottom if the degree on top is higher than the degree on bottom like this one we had up here then there is no horizontal asymptote if the degree on top is lower than the degree on bottom then the horizontal asymptote is y=0

OpenStudy (anonymous):

PERFECT! :D...thanks!!!

myininaya (myininaya):

so we have (3x^2+1)/(6x^2-9) then horizontal asymptote is y=1/2

myininaya (myininaya):

if we have that is an example i made up k?

myininaya (myininaya):

so if we have (x-1)^2/(x+1)^3 then the horizontal asymptote is y=0

myininaya (myininaya):

what do you think the horiztontal asymptote is of (2x-1)^3/(x-1)^2

myininaya (myininaya):

degree on top>degree on bottom so there is not one

OpenStudy (anonymous):

i see

myininaya (myininaya):

(2x-1)^3/(x-1)^3 this one?

OpenStudy (anonymous):

first one...both have same exponents, so the asymptote is their coefficients

myininaya (myininaya):

my first example to you yes i said y=1/2 or you could have said y=3/6

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

top degree > bottom degree = no asymptote top degree < bottom degree = asymptote = 0 equal degree on denominator and numerator = leading coefficients

myininaya (myininaya):

so my last example the degrees are the same right? so on top you have after expanding it you have 8x^3 + gibberish on bottom you have after expanding it you have x^3 + gibberish so the horizontal asymptote is y=8

OpenStudy (anonymous):

thanks! :D

myininaya (myininaya):

right

myininaya (myininaya):

leading coefficient of top --------------------- leading coefficicent of bottom if the degree of top=degree of bottom

OpenStudy (anonymous):

yes

myininaya (myininaya):

:)

OpenStudy (anonymous):

very very helpful!